# On the Validity of the Background Subtraction Method for Black Hole Thermodynamics in Matter–Coupled Gravity Theories

### v1

@@ I Introduction
On the other hand, the Iyer–Wald formalism is another powerful method for studying black hole thermodynamics Wald:1993nt ; Iyer:1994ys ; Iyer:1995kg . These early works have already explained that the results obtained from the Euclidean method are consistent with the Noether–charge construction. Although the essential idea was known long ago, a refined and fully unambiguous explanation of how background subtraction emerges naturally from the Iyer–Wald formalism has appeared only recently in Guo:2025ohn ; Chen:2025ary .
@@ II The derivation of background subtraction method from Iyer–Wald formalism
Based on the excellent earlier developments Wald:1993nt ; Iyer:1994ys ; Iyer:1995kg ; Guo:2025ohn ; Chen:2025ary , we now have a clear understanding of how the background subtraction method can be naturally induced from the Iyer–Wald formalism, and why it is expected to be valid.
@@
The left-hand side of Eq. (16) is known to equal $T\,\delta S$, where $T$ is the Hawking temperature and $S$ is the black hole entropy Wald:1993nt ; Iyer:1994ys . For the right-hand side, we can evaluate $\bm{\Theta}$ on the timelike boundary $\Gamma$ at infinity, where $S_{\infty}=\Sigma\cap\Gamma$ and the induced volume form $\hat{\bm{\epsilon}}$ on $\Gamma$ satisfies $\bm{\epsilon}=\bm{n}\wedge\hat{\bm{\epsilon}}$ (where $\bm{n}$ being the unit normal vector to $\Gamma$). Following standard procedures Padmanabhan:2014lwa ; Chakraborty:2017zep ; Jiang:2018sqj , $\bm{\Theta}$ at the boundary can be decomposed as
@@
Here, $K$ is the extrinsic curvature of $\Gamma$, and $h_{\mu\nu}$ is the induced metric. The $d\bm{C}$ term has a well-defined expression but vanishes in calculations, so we omit it throughout this work  [^2] While the $d\bm{C}$ term vanishes numerically, it is useful for formal derivations, e.g., rewriting Eq. (26) in the Brown-York mass form. See Harlow:2019yfa ; Guo:2024oey for details..
@@
Next, we demonstrate how the quantum statistical relation $I_{E}=\beta(M-TS)$ arises naturally, where $\beta\equiv 1/T$ is the inverse of Hawking temperature and the periodicity of the Euclidean time $\tau$. The derivation is as follows:
@@
In the first step, we use the definition of $M$ in Eq. (26) and the key result from the Iyer–Wald formalism: $\int^{(\mathrm{BH})}_{S_{h}}\bm{Q}=TS_{W}$, where $S_{W}$ denotes the well-known Wald entropy Iyer:1994ys ; Iyer:1995kg . The second step invokes the integral form of Eq. (9), which means $\oint_{S}\bm{Q}=-\int_{\Sigma}\xi_{t}\cdot\bm{L}$. For the third step, we simply abbreviate the expressions using the regularized integral $\int^{(\mathrm{Reg.})}\equiv\int^{(\mathrm{BH})}-\int^{(\mathrm{bg})}$. In the final step, we perform a Wick rotation $t=-i\tau$, with $\bm{L}_{E}=-i\bm{L}$ and $\bm{B}_{E}=-i\bm{B}$. Specifically, we have $\beta\int\xi_{t}\cdot\bm{L}=\int d\tau\wedge\frac{\partial}{\partial\tau}\cdot\bm{L}_{E}=\int\bm{L}_{E}$. A similar derivation applies to $\bm{B}_{E}$, with an additional sign flip caused by the orientation convention of $\partial\mathcal{M}$.
@@ III.1 Higher-derivative gravity with non-minimally coupled Maxwell field
$\bm{L}=\big(R-2\Lambda-\frac{1}{4}F^{2}+\gamma\mathcal{L}_{\text{hd}}\big)\bm{\epsilon},$ (31)
@@
$Q$ $=\frac{\partial G}{\partial\Phi}=\frac{q}{4},$ (56)
@@ III.2 Gravity coupled with a background Kalb-Ramond field
However, we now present a case that serves as a cautionary note: when moving beyond pure gravity and AdS spacetimes, Eq. (59) may fail to yield correct thermodynamic quantities. This occurs when matter fields possess non-vanishing vacuum expectation values that approach finite constants at infinity, leading to nonzero boundary term contributions. We illustrate this with gravity coupled to a Kalb-Ramond field, which is a second-rank antisymmetric tensor field from the bosonic spectrum of string theory. As far as we know, Euclidean action computations for such black holes have not been reported previously.
@@ Acknowledgments
YX would like to thank Hong-Bao Zhang, Xi-Yao Guo, and Jie-Qiang Wu for useful discussions. YX is also thankful to the Higgs Centre for Theoretical Physics at the University of Edinburgh for providing research facilities and hospitality during the visit. This work was supported in part by China Scholarship Council with Grant No. 202408130101, the Hebei Natural Science Foundation with Grant No. A2024201012, and the National Natural Science Foundation of China with Grant No. 12475048.

### v2

@@ I Introduction
On the other hand, the Iyer–Wald formalism is another powerful method for studying black hole thermodynamics Wald:1993nt ; Iyer:1994ys ; Iyer:1995kg . These early works have already explained that the results obtained from the Euclidean method are consistent with the Noether–charge construction. Although the essential idea was known long ago, a refined and fully unambiguous explanation of how background subtraction emerges naturally from the Iyer–Wald formalism has appeared only recently in Guo:2025ohn ; Guo:2025muo ; Chen:2025ary .
@@ II The derivation of background subtraction method from Iyer–Wald formalism
Based on the excellent earlier developments Wald:1993nt ; Iyer:1994ys ; Iyer:1995kg ; Guo:2025ohn ; Guo:2025muo ; Chen:2025ary , we now have a clear understanding of how the background subtraction method can be naturally induced from the Iyer–Wald formalism, and why it is expected to be valid.
@@
The left-hand side of Eq. (16) is known to equal $T\,\delta S$, where $T$ is the Hawking temperature and $S$ is the black hole entropy Wald:1993nt ; Iyer:1994ys . For the right-hand side, we can evaluate $\bm{\Theta}$ on the timelike boundary $\Gamma$ at infinity, where $S_{\infty}=\Sigma\cap\Gamma$ and the induced volume form $\hat{\bm{\epsilon}}$ on $\Gamma$ satisfies $\bm{\epsilon}=\bm{n}\wedge\hat{\bm{\epsilon}}$ (where $\bm{n}\propto\nabla r$ being the outward unit normal vector to $\Gamma$). Following standard procedures Padmanabhan:2014lwa ; Chakraborty:2017zep ; Jiang:2018sqj , $\bm{\Theta}$ at the boundary can be decomposed as
@@
Here, $K$ is the extrinsic curvature of $\Gamma$, and $h_{\mu\nu}$ is the induced metric. The $d\bm{C}$ term also has a well-defined expression but vanishes in calculations, so we omit it throughout this work  [^2] While the $d\bm{C}$ term vanishes numerically, it is useful for formal derivations, e.g., rewriting Eq. (26) in the Brown-York mass form. See Harlow:2019yfa ; Guo:2024oey for details..
@@
Next, we demonstrate how the quantum statistical relation $I_{E}=\beta(M-TS_{W})$ arises naturally, where $\beta\equiv 1/T$ is the inverse of Hawking temperature and the periodicity of the Euclidean time $\tau$. The derivation is as follows:
@@
In the first step, we use the definition of $M$ in Eq. (26) and the key result from the Iyer–Wald formalism: $\int^{(\mathrm{BH})}_{S_{h}}\bm{Q}=TS_{W}$, where $S_{W}$ denotes the well-known Wald entropy Iyer:1994ys ; Iyer:1995kg . The second step invokes the integral form of Eq. (9), which means $\oint_{S}\bm{Q}=-\int_{\Sigma}\xi_{t}\cdot\bm{L}$. For the third step, we simply abbreviate the expressions using the regularized integral $\int^{(\mathrm{Reg.})}\equiv\int^{(\mathrm{BH})}-\int^{(\mathrm{bg})}$. In the final step, we perform a Wick rotation $t=-i\tau$, with $\bm{L}_{E}=-i\bm{L}$ and $\bm{B}_{E}=-i\bm{B}$. Specifically, we have $\beta\int\xi_{t}\cdot\bm{L}=\int d\tau\wedge\frac{\partial}{\partial\tau}\cdot\bm{L}_{E}=\int\bm{L}_{E}$. A similar derivation applies to $\bm{B}_{E}$, with an additional sign flip caused by the orientation convention of the boundary $\Gamma$ (denoted as $\partial\mathcal{M}$ as well).
@@ III.1 Higher-derivative gravity with non-minimally coupled Maxwell field
$\bm{L}=\frac{1}{16\pi}\big(R-2\Lambda-\frac{1}{4}F^{2}+\gamma\mathcal{L}_{\text{hd}}\big)\bm{\epsilon},$ (31)
@@
$Q$ $=-\frac{\partial G}{\partial\Phi}=\frac{q}{4},$ (56)
@@ III.2 Gravity coupled with a background Kalb-Ramond field
However, we now present a case that serves as a cautionary note: when moving beyond the regime of pure gravity or AdS spacetimes, Eq. (59) may fail to yield correct thermodynamic quantities. This occurs when matter fields possess non-vanishing vacuum expectation values that approach finite constants at infinity, leading to nonzero boundary term contributions. We illustrate this with gravity coupled to a Kalb-Ramond field, which is a second-rank antisymmetric tensor field from the bosonic spectrum of string theory. As far as we know, Euclidean action computations for such black holes have not been reported previously.
@@ Acknowledgments
YX would like to thank Hong-Bao Zhang, Xi-Yao Guo and Jie-Qiang Wu for useful discussions. YX is also thankful to the Higgs Centre for Theoretical Physics at the University of Edinburgh for providing research facilities and hospitality during the visit. This work was supported in part by China Scholarship Council with Grant No. 202408130101, the Hebei Natural Science Foundation with Grant No. A2024201012, and the National Natural Science Foundation of China with Grant No. 12475048.